# Hybrid estimation for a mixed fractional Black-Scholes model with random effects from discrete time observations

### v3

@@ 7.11 Proof of Proposition 7
For all $x\in[-1,1]$, we decompose the error of $\hat{F}_{m,n,N}$ as follows
$\left|\hat{F}_{m,n,N}(x)-F(x)\right|\leq\left|\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right|+\left|\hat{F}_{N}(x)-\bar{F}_{N}(x)\right|+\left|\bar{F}_{N}(x)-F(x)\right|,$ (7.36)
where $\hat{F}_{N}(x)=\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{\hat{\phi}_{i}\leq x\}}$ is the empirical distribution function of the random effects estimators $\hat{\phi}_{i}$ and $\bar{F}_{N}(x)=\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{\phi_{i}\leq x\}}$ is the empirical distribution function of the unknown random effects $\phi_{i}$. For the first term of the right hand side of (7.36), from (7.31), there exists a positive constant $C$ such that
$\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right|\leq Cm^{-2}.$
Multiplying by the scaling factor, we get
$\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right|\leq C\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}m^{-2}.$
If we choose $m=o\left(\left(\frac{\log\log N}{N}\right)^{\frac{1}{4}}\right)$, then $\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}m^{-2}\to 0$, as $N\to\infty$, which implies that
$\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{m,n,N}(x)-\hat{F}_{N}(x)\right|\to 0,\quad\text{as}\quad N\to\infty.$ (7.37)
We now analyze the second term of (7.36). For all $x\in[-1,1]$, we have
$\left|\hat{F}_{N}(x)-\bar{F}_{N}(x)\right|\leq\dfrac{1}{N}\sum\limits_{i=1}^{N}\left|\mathds{1}_{\{\hat{\phi}_{i}\leq x\}}-\mathds{1}_{\{\phi_{i}\leq x\}}\right|\leq\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{|\hat{\phi}_{i}-\phi_{i}|>\varepsilon\}},\quad\text{for any}\ \varepsilon>0.$
Hence,
$\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{N}(x)-\bar{F}_{N}(x)\right|\leq\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{|\hat{\phi}_{i}-\phi_{i}|>\varepsilon\}}.$
Since for all $i=1,\ldots,N$, $\hat{\phi}_{i}\xrightarrow{a.s}\phi_{i}$ as $n,N\to\infty$, the SLLN implies that
$\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{|\hat{\phi}_{i}-\phi_{i}|>\varepsilon\}}\xrightarrow{a.s}\mathbb{P}\left(|\hat{\phi}_{i}-\phi_{i}|>\varepsilon\right)=0,\quad\text{as}\ \ n,N\to\infty.$
Furthermore, since $\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}\frac{1}{N}=o(1)$, as $N\to\infty$, it follows that
$\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\dfrac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{|\hat{\phi}_{i}-\phi_{i}|>\varepsilon\}}\xrightarrow{a.s}0,\quad\text{a.s}\ n,N\to\infty.$
Therefore,
$\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{N}(x)-\bar{F}_{N}(x)\right|\xrightarrow{a.s}0,\quad\text{as}\ \ n,N\to\infty.$ (7.38)
Now, we proceed to analyze the last term of (7.36). Since $\bar{F}_{n}$ is the empirical distribution function of the i.i.d random effects $\phi_{i}$, it is well known that it verifies the Chung-Smirnov property, that is
$\underset{N\to\infty}{\lim\sup}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\left|\bar{F}_{N}(x)-F(x)\right|=1,\ \text{a.s}.$ (7.39)
Finally, combining (7.37), (7.38) and (7.39), we deduce
$\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{m,n,N}(x)-F(x)\right|\leq o(1)+o(1)+\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\bar{F}_{N}(x)-F(x)\right|,$
Taking the $\lim\sup$ when NN goes to infinity, we obtain
$\underset{N\to\infty}{\lim\sup}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\underset{x\in[-1,1]}{\sup}\left|\hat{F}_{m,n,N}(x)-F(x)\right|\leq 1,$
@@ Appendix A Appendix
Let $\{X_{i},\ i\geq 1\}$ be a stationary ergodic sequence of $\mathbb{R}^{k}$ valued random variables on some probability space $\Big(\Omega,\mathcal{F},\mathbb{F},\mathbb{P}\Big)$ and let $g:\mathbb{R}^{k}\longrightarrow\mathbb{R}$ be a Borel measurable function such that $\mathbb{E}\left(\mid g\left(X_{1},X_{2},\ldots,X_{k}\right)\mid\right)<\infty$. Then

### v4

@@ 7.11 Proof of Proposition 7
For all $x\in[-1,1]$, we decompose the error as
$\left|\widehat{F}_{m,n,N}(x)-F(x)\right|\leq\left|\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\right|+\left|\widehat{F}_{N}(x)-\bar{F}_{N}(x)\right|+\left|\bar{F}_{N}(x)-F(x)\right|,$ (7.47)
where
$\widehat{F}_{N}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{\widehat{\phi}_{i}\leq x\}},\qquad\bar{F}_{N}(x)=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{\phi_{i}\leq x\}}.$
We begin with the first term of (7.47). From (7.37), there exists $C^{\prime}>0$ such that
$\sup_{x\in[-1,1]}\left|\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\right|\leq C^{\prime}\,\frac{\log m}{m^{2}}.$
Multiplying by the Chung–Smirnov scaling factor yields
$\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\sup_{x\in[-1,1]}\left|\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\right|\leq C^{\prime}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\frac{\log m}{m^{2}}.$ (7.48)
Taking $m=m_{N}$ that verifies $\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\frac{\log m}{m^{2}}\to 0$ as $N\to\infty$, the right-hand side converges to zero, and therefore
$\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}\sup_{x\in[-1,1]}\left|\widehat{F}_{m,n,N}(x)-\widehat{F}_{N}(x)\right|\longrightarrow 0,\qquad\text{a.s.}$ (7.49)
For the second term of the decomposition (7.47), as we have shown in (7.36),
$\sup_{x\in[-1,1]}|\widehat{F}_{N}(x)-\bar{F}_{N}(x)|\xrightarrow{a.s.}0,$ (7.50)
Hence, multiplying by the Chung-Smirnov scaling factor, we obtain
$\left(\frac{2N}{\log\log N}\right)^{\frac{1}{2}}\sup_{x\in[-1,1]}|\widehat{F}_{N}(x)-\bar{F}_{N}(x)|\xrightarrow{a.s.}0.$
For the last term of (7.47), since $\bar{F}_{N}$ is the empirical distribution function of i.i.d. variables $\phi_{i}$, it satisfies
$\limsup_{N\to\infty}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\sup_{x\in[-1,1]}\left|\bar{F}_{N}(x)-F(x)\right|=1,\qquad\text{a.s.}$ (7.51)
Finally, by Combining (7.49), (7.50), and (7.51) in (7.47), we conclude that
$\limsup_{N\to\infty}\left(\dfrac{2N}{\log\log N}\right)^{\frac{1}{2}}\sup_{x\in[-1,1]}\left|\widehat{F}_{m,n,N}(x)-F(x)\right|\leq 1,\qquad\text{a.s.},$
@@
7.12 Proof of Theorem 4
Throughout this proof, we write $m=m_{N}$, $n=n_{N}$. By the triangle inequality, the squared error of $\widehat{F}_{m,n,N}$ can be decomposed as follows
$\|\widehat{F}_{m,n,N}-F\|^{2}\leq 3\Bigl(\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|^{2}+\|\widehat{F}_{N}-\bar{F}_{N}\|^{2}+\|\bar{F}_{N}-F\|^{2}\Bigr).$ (7.52)
For the interpolation error, we have from (7.37) in the proof of Theorem 2,
$\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|\leq C^{\prime}\frac{\log m}{m^{2}},$
which yields
$\mathbb{E}\bigl[\|\widehat{F}_{m,n,N}-\widehat{F}_{N}\|^{2}\bigr]\leq C^{\prime 2}\frac{(\log m)^{2}}{m^{4}}.$ (7.53)
For the last term of (7.52), since the random variables $\phi_{i}$, $i=1,\ldots,N$, are i.i.d with distribution function $F$, the Dvoretzky–Kiefer–Wolfowitz (DKW) inequality ensures that for every $s>0$,
$\mathbb{P}\bigl(\|\bar{F}_{N}-F\|>s\bigr)\leq 2e^{-2Ns^{2}}.$
Since $\|\bar{F}_{N}-F\|$ is a non-negative random variable, its expectation verifies
$\mathbb{E}\bigl[\|\bar{F}_{N}-F\|^{2}\bigr]$ $=\int_{0}^{\infty}2s\,\mathbb{P}\bigl(\|\bar{F}_{N}-F\|>s\bigr)ds$
$\leq\int_{0}^{\infty}2s\,\min\,\!\bigl(1,2e^{-2Ns^{2}}\bigr)ds.$
The last integral can be split into two parts based on the value of $s_{0}$ that verifies $2e^{-2Ns_{0}^{2}}=1$, i.e. $s_{0}=\Big(\frac{\log 2}{2N}\Big)^{\frac{1}{2}}$. Hence,
$\mathbb{E}\bigl[\|\bar{F}_{N}-F\|^{2}\bigr]\leq\frac{1+\frac{\log 2}{2}}{N}\leq\frac{2}{N}.$ (7.54)
We now proceed to find an upper bound for the second term of (7.52). Using (7.32), we can fix a sequence $\varepsilon_{N}\to 0$ such that, along the chosen diagonal $n=n_{N}$,
$\max\limits_{1\leq i\leq N}|\widehat{\phi}_{i}-\phi_{i}|\leq\varepsilon_{N}\quad\text{a.s}\quad\text{for all sufficiently large }N.$
Moreover, from (7.30), we have for any $x\in[-1,1]$,
$\|\widehat{F}_{N}-\bar{F}_{N}\|\leq\frac{1}{N}\sum\limits_{i=1}^{N}\mathds{1}_{\{|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon_{N}\}}+\sup_{x\in[-1,1]}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon_{N}<\phi_{i}\leq x+\varepsilon_{N}\}}\quad\text{a.s.\ for large }N.$ (7.55)
Squaring and using $(a+b)^{2}\leq 2(a^{2}+b^{2})$,
$\|\widehat{F}_{N}-\bar{F}_{N}\|^{2}\leq 2\Bigl(\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon_{N}\}}\Bigr)^{2}+2\Bigl(\sup\limits_{x\in[-1,1]}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon_{N}<\phi_{i}\leq x+\varepsilon_{N}\}}\Bigr)^{2}.$
For $N$ large enough, we have
$\mathbb{E}\Bigl[\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{|\widehat{\phi}_{i}-\phi_{i}|>\varepsilon_{N}\}}^{2}\Bigr]=0.$
For the second term, since $\phi_{i}$ have bounded density $f$ under assumption $(A3)$,
$\mathbb{E}\Bigl[\sup\limits_{x\in[-1,1]}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon_{N}<\phi_{i}\leq x+\varepsilon_{N}\}}\Bigr]\leq 2\|f\|\varepsilon_{N},$
and by the uniform law of large numbers
$\mathrm{Var}\Bigl(\sup_{x}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon<\phi_{i}\leq x+\varepsilon\}}\Bigr)=O\!\left(\frac{\varepsilon}{N}\right).$
Hence
$\mathbb{E}\Bigl[\Bigl(\sup\limits_{x\in[-1,1]}\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{\{x-\varepsilon_{N}<\phi_{i}\leq x+\varepsilon_{N}\}}\Bigr)^{2}\Bigr]\leq 4\|f\|^{2}\varepsilon_{N}^{2}+O\!\left(\frac{\varepsilon_{N}}{N}\right).$
Replacing into (7.55) and squaring, we obtain
$\mathbb{E}\bigl[\|\widehat{F}_{N}-\bar{F}_{N}\|^{2}\bigr]\leq C_{f}\,\varepsilon_{N}^{2}+o\left(\frac{\varepsilon_{N}}{N}\right),\qquad C_{f}:=4\|f\|^{2}.$ (7.56)
Finally, inserting (7.54), (7.53) and (7.56) into (7.52), we obtain the needed bound.
@@ Appendix A Appendix
Let $\{X_{i},\ i\geq 1\}$ be a stationary ergodic sequence of $\mathbb{R}^{k}$ valued random variables on some probability space $\Big(\Omega,\mathcal{F},\mathbb{F},\mathbb{P}\Big)$ and let $g:\mathbb{R}^{k}\longrightarrow\mathbb{R}$ be a Borel measurable function such that
$\mathbb{E}\left(\mid g\left(X_{1},X_{2},\ldots,X_{k}\right)\mid\right)<\infty$. Then